OpenStudy (anonymous):

how do you distribute x^-1/2

OpenStudy (anonymous):

x/1/2=2x

OpenStudy (anonymous):

ok how about one step further. \[\frac{ 1 }{ 2 }(5-3x)^{-1/2}\times (-3)\]

OpenStudy (anonymous):

can you show me step process?

OpenStudy (anonymous):

well i just did it simply =)

OpenStudy (anonymous):

whats the answer?

OpenStudy (anonymous):

okay let me show u -- x^-1/2 =x/1/2=x.2/1=2x/1=2x

OpenStudy (anonymous):

no I'm asking how I can solve the equation I wrote just now :) thanks about 2x though

OpenStudy (anonymous):

1/2.(5-3x)^-1/2.(-3) =1/2(5-3x/1/2).(-3) =1/2.2(5-3x).(-3) =10-6x/2.(-3) =2(5-3x)/2.(-3) =5-3x.(-3) =-15+9x =9x-15 =3(3x-5)

OpenStudy (anonymous):

thanks!